OpenStudy (anonymous):

What is the product of √2 (6b√16 - 2b)? Simplify if possible.

OpenStudy (anonymous):

16b 4b√2 22b√2 16b√2

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

i don't understand... sorry its been a while since i've done this trying to help my son